# Improving Drone Racing Performance Through Iterative Learning MPC

### v2

@@ III Problem Definition
From a fixed starting pose $\mathbf{T}_{is}=\{\mathbf{p}^{si}_{i},\mathbf{R}_{is}\}\in SE(3)$ (where $i$ indicates the inertial frame $I$ and $s$ indicates the starting frame $S$), the drone is tasked to pass through a series of $N_{g}$ gates. Each gate is defined by a given convex and compact set $\mathbb{T}_{n}\subset\mathbb{R}^{3}$ such that $\mathbf{t}^{\intercal}\mathbf{e}_{z}=0\,,\forall\mathbf{t}\in\mathbb{T}_{n}$ with $\mathbf{e}_{z}=\left[0\quad 0\quad 1\right]^{\intercal}$ indicating the opening of the gate and a known pose in the inertial frame $\mathbf{T}_{it_{n}}\in SE(3)$. Each gate’s $z$-direction indicates the direction of traversal, and the drone must traverse all gates in the given order $n\in\{1,\dots,N_{g}\}$. We refer to this setup as the track. The task is considered to be executed successfully if no constraints are violated (e.g., no collisions with the gates or the environment). Since we consider the racing setting, the goal is to minimize the time required for the drone to complete the track successfully. The traversal time is the time $t_{N_{g}}$ indicating when the drone passes through the last gate with index $N_{g}$ after starting from the initial pose at time $t_{0}=0$. Mathematically, this can be formulated as a minimum-time optimal control problem:
@@ IV-A Drone Dynamics
The drone’s state is defined as $\mathbf{x}=[\mathbf{p}_{i}^{\intercal},\mathbf{v}_{i}^{\intercal},\bm{\phi}^{ib\intercal}_{i}]^{\intercal}\in\mathbb{R}^{9}$, where $\mathbf{p}_{i}\in\mathbb{R}^{3}$ is the position of the drone’s center of gravity (CoG) given in the inertial frame $I$, $\mathbf{v}_{i}\in\mathbb{R}^{3}$ is the velocity, $\bm{\phi}^{ib}_{i}\in SO(3)$ are the Euler angles that describe the rotation from the body frame $B$ to the inertal frame. The control input is the collective thrust-attitude interface $\mathbf{u}=[f_{\Sigma},\bm{\phi}_{\text{cmd}}^{\intercal}]^{\intercal}\in\mathbb{R}^{4}$, where $f_{\Sigma}=\sum_{i=1}^{4}f_{i}\in\mathbb{U}_{f}$ is the collective thrust from the four propellers, $\mathbb{U}_{f}=\left[0,f_{\Sigma,\max}\right]$, and $\bm{\phi}_{\text{cmd}}\in SO(3)$ are the commanded Euler angles. Therefore, $\mathbb{U}=\mathbb{U}_{f}\times SO(3)$. For brevity, we drop the frame indices as all the variables are expressed in the inertial frame, which leads to the following state and input vector expressions: $\mathbf{x}=[x,y,z,v_{x},v_{y},v_{z},\phi,\theta,\psi]^{\intercal}\,,\text{ and}~\mathbf{u}=[f_{\Sigma},\phi_{\text{cmd}},\theta_{\text{cmd}},\psi_{\text{cmd}}]^{\intercal}\,.$ Then, the translational dynamics can be derived from Newton’s law, and the rotational dynamics can be modeled as three independent first-order integrators:
$\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})+\begin{bmatrix}\mathbf{G}_{1}(\mathbf{x})&\mathbf{0}\\ \mathbf{0}&\mathbf{G}_{2}\end{bmatrix}\begin{bmatrix}f_{\Sigma}\\ \bm{\phi}_{\text{cmd}}\end{bmatrix}\,$, (3)
@@
$\mathbf{G}_{1}(\mathbf{x})=\frac{1}{m}\begin{bmatrix}\text{c}\phi\,\text{s}\theta\,\text{c}\psi+\text{s}\phi\,\text{s}\psi\\ \text{c}\phi\,\text{s}\theta\,\text{s}\psi-\text{s}\phi\,\text{c}\psi\\ \text{c}\phi\,\text{c}\theta\end{bmatrix}f_{\Sigma}\,,~\mathbf{G}_{2}=\text{diag}(\beta_{\phi}\,,\beta_{\theta}\,,\beta_{\psi})\,$,
@@ VI-B Solver Runtime Statistics
To ensure real-time feasibility of our proposed LMPC variant, we analyze its computational cost under different hyperparameter configurations. Compared to the vanilla LMPC, our proposed variant introduces an additional runtime module for arc-length estimation. For a global safe set consisting of 631 states, directly computing $s$ via brute-force linear search takes on average $4.21\pm 0.28$ ms per call. To reduce this overhead, we employ a $k$-d tree structure, which reduces the estimation to $0.68\pm 0.17$ ms and is negligible compared to the overall solver time. Therefore, we always use the $k$-d tree in our implementations. Table I reports the average runtime of the proposed LMPC variant as a function of the prediction horizon $N$ and the size of the safe set $K$. Based on this trade-off between computation time and performance, we select $(N,K)=(8,20)$ as the default setting for the following simulation and real-world experiments. This setting offers a good balance between long-term prediction and optimization feasibility, while maintaining a control frequency of 30 Hz.
@@ VI-C1 Ablation Study on LMPC Hyperparameters
The results (see Figure 4) indicate that LMPC iterations primarily lead to two effects: acceleration and shortcutting. The acceleration effect is primarily driven by a smaller discretization frequency $f_{d}$, while the influence of $K$ is relatively minor. In contrast, shortcutting behavior is promoted by both a smaller $f_{d}$ and a larger safe set size $K$. Unlike acceleration, which is purely beneficial, shortcutting can be detrimental in drone racing. Due to the existence of model mismatch and state estimation errors, overly aggressive shortcuts increase the risk of collisions. Therefore, in the context of drone racing, choosing relatively small $f_{d}$ and $K$ with an appropriate $N$ yields better overall performance and safety.

### v3

@@ III Problem Definition
From a fixed starting pose $\mathbf{T}_{is}=\{\mathbf{p}^{si}_{i},\mathbf{R}_{is}\}\in SE(3)$ (where $i$ indicates the inertial frame II and $s$ indicates the starting frame $S$), the drone is tasked to pass through a series of $N_{g}$ gates. Each gate is defined by a given convex and compact set $\mathbb{T}_{n}\subset\mathbb{R}^{3}$ such that $\mathbf{t}^{\intercal}\mathbf{e}_{z}=0\,,\forall\mathbf{t}\in\mathbb{T}_{n}$ with $\mathbf{e}_{z}=\left[0\quad 0\quad 1\right]^{\intercal}$ indicating the opening of the gate and a known pose in the inertial frame $\mathbf{T}_{it_{n}}\in SE(3)$. Each gate’s $z$-direction indicates the direction of traversal, and the drone must traverse all gates in the given order $n\in\{1,\dots,N_{g}\}$. We refer to this setup as the track. The task is considered to be executed successfully if no constraints are violated (e.g., no collisions with the gates or the environment). Since we consider the racing setting, the goal is to minimize the time required for the drone to complete the track successfully. The traversal time is the time $t_{N_{g}}$ indicating when the drone passes through the last gate with index $N_{g}$ after starting from the initial pose at time $t_{0}=0$. Mathematically, this can be formulated as a minimum-time optimal control problem:
@@ IV-A Drone Dynamics
The drone’s state is defined as $\mathbf{x}=[\mathbf{p}_{i}^{\intercal},\mathbf{v}_{i}^{\intercal},\bm{\phi}^{ib\intercal}_{i}]^{\intercal}\in\mathbb{R}^{9}$, where $\mathbf{p}_{i}\in\mathbb{R}^{3}$ is the position of the drone’s center of gravity (CoG) given in the inertial frame II, $\mathbf{v}_{i}\in\mathbb{R}^{3}$ is the velocity, $\bm{\phi}^{ib}_{i}\in SO(3)$ are the Euler angles that describe the rotation from the body frame $B$ to the inertal frame. The control input is the collective thrust-attitude interface $\mathbf{u}=[f_{\Sigma},\bm{\phi}_{\text{cmd}}^{\intercal}]^{\intercal}\in\mathbb{R}^{4}$, where $f_{\Sigma}=\sum_{i=1}^{4}f_{i}\in\mathbb{U}_{f}$ is the collective thrust from the four propellers, $\mathbb{U}_{f}=\left[0,f_{\Sigma,\max}\right]$, and $\bm{\phi}_{\text{cmd}}\in SO(3)$ are the commanded Euler angles. Therefore, $\mathbb{U}=\mathbb{U}_{f}\times SO(3)$. For brevity, we drop the frame indices as all the variables are expressed in the inertial frame, which leads to the following state and input vector expressions: $\mathbf{x}=[x,y,z,v_{x},v_{y},v_{z},\phi,\theta,\psi]^{\intercal}\,,\text{ and}~\mathbf{u}=[f_{\Sigma},\phi_{\text{cmd}},\theta_{\text{cmd}},\psi_{\text{cmd}}]^{\intercal}\,.$ Then, the translational dynamics can be derived from Newton’s law, and the rotational dynamics can be modeled as three independent first-order integrators:
$\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})+\begin{bmatrix}\mathbf{0}_{3\times 1}&\mathbf{0}_{3\times 3}\\ \mathbf{G}_{1}(\mathbf{x})&\mathbf{0}_{3\times 3}\\ \mathbf{0}_{3\times 1}&\mathbf{G}_{2}\end{bmatrix}\begin{bmatrix}f_{\Sigma}\\ \bm{\phi}_{\text{cmd}}\end{bmatrix}\,$, (3)
@@
$\mathbf{G}_{1}(\mathbf{x})=\frac{1}{m}\begin{bmatrix}\text{c}\phi\,\text{s}\theta\,\text{c}\psi+\text{s}\phi\,\text{s}\psi\\ \text{c}\phi\,\text{s}\theta\,\text{s}\psi-\text{s}\phi\,\text{c}\psi\\ \text{c}\phi\,\text{c}\theta\end{bmatrix}\,,~\mathbf{G}_{2}=\text{diag}(\beta_{\phi}\,,\beta_{\theta}\,,\beta_{\psi})\,$,
@@ VI-B Solver Runtime Statistics
To ensure real-time feasibility of our proposed LMPC variant, we analyze its computational cost under different hyperparameter configurations. Compared to the vanilla LMPC, our proposed variant introduces an additional runtime module for arc-length estimation. For a global safe set consisting of 631 states, directly computing $s$ via brute-force linear search takes on average $4.21\pm 0.28$ ms per call. To reduce this overhead, we employ a $k$-d tree structure, which reduces the estimation to $0.68\pm 0.17$ ms and is negligible compared to the overall solver time. Therefore, we always use the $k$-d tree in our implementations. Table I reports the average runtime of the proposed LMPC variant as a function of the prediction horizon NN and the size of the safe set $K$. Based on this trade-off between computation time and performance, we select $(N,K)=(8,20)$ as the default setting for the following simulation and real-world experiments. This setting offers a good balance between long-term prediction and optimization feasibility, while maintaining a control frequency of 30 Hz.
@@ VI-C1 Ablation Study on LMPC Hyperparameters
The results (see Figure 4) indicate that LMPC iterations primarily lead to two effects: acceleration and shortcutting. The acceleration effect is primarily driven by a smaller discretization frequency $f_{d}$, while the influence of $K$ is relatively minor. In contrast, shortcutting behavior is promoted by both a smaller $f_{d}$ and a larger safe set size $K$. Unlike acceleration, which is purely beneficial, shortcutting can be detrimental in drone racing. Due to the existence of model mismatch and state estimation errors, overly aggressive shortcuts increase the risk of collisions. Therefore, in the context of drone racing, choosing relatively small $f_{d}$ and $K$ with an appropriate NN yields better overall performance and safety.